### v1

# Large Language Model Aided QoS Prediction for Service Recommendation

Abstract
Large language models (LLMs) have seen rapid improvement in the recent years, and are used in a wider range of applications. After being trained on large text corpus, LLMs obtain the capability of extracting rich features from textual data. Such capability is potentially useful for the web service recommendation task, where the web users and services have intrinsic attributes that can be described using natural language sentences and are useful for recommendation. In this paper, we explore the possibility and practicality of using LLMs for web service recommendation. We propose the large language model aided QoS prediction (llmQoS) model, which use LLMs to extract useful information from attributes of web users and services via descriptive sentences. This information is then used in combination with the QoS values of historical interactions of users and services, to predict QoS values for any given user-service pair. Our proposed model is shown to overcome the data sparsity issue for QoS prediction. We show that on the WSDream dataset, llmQoS outperforms comparable baseline models consistently.
Index Terms:
Service Recommendation, Quality of Service, Large Language Model.
I Introduction
FIGURE I: Illustration of the QoS prediction problem. A QoS matrix exists for sets of web users and services, and each elements of the matrix corresponds to a user-service interaction. A QoS prediction model aims to predict the value of unknown QoS values from the known QoS values from historical user-service interactions.
The rapid development of cloud computing has led to an explosive growth in the number of cloud services, resulting in many services having identical or similar functionalities [1]. There is a growing need for service providers to provide the most suitable service for users to meet customer demands. Where there are many candidate services for a user request, quality of service (QoS) is a commonly used metric to measure the performance of these services. QoS refers to non-functionality aspects of services [2] such as throughput, response time, and cost. It plays a crucial role in distinguishing similar services and helps users select the most appropriate one. However, in real-world scenarios, evaluating the QoS of a large number of services on the client side is expensive and time-consuming due to the sheer scale. Additionally, most end users lack the expertise to do the evaluation [3]. Therefore, the only feasible option for an end user is to predict the QoS values for an unknown service, and the QoS values are used to help the user find the most suitable service.
The problem of QoS prediction is illustrated in Figure I. For the sets of web users and services exists a QoS matrix, in which each value corresponds to a user-service interaction. The matrix contains some known QoS values from historical user-service interactions, and our goal is to predict unknown QoS values using these known values, thereby recommending more suitable services to users. Many methods have been developed to predict the QoS values for a given user and service. Among them, collaborative filtering is the most widely used technique. It aim to accurately estimate QoS values by learning from the QoS value matrix from historical user-service interactions. However, in practice, the historical QoS value matrix is often highly sparse, as each user usually only interacted with a limited number of cloud services in the past. This data sparsity issue poses difficulty for many QoS prediction methods. Although many methods aim to mitigate this difficulty, it is still challenging to achieve ideal predictive performance.
More recently, large language models (LLMs) have demonstrated remarkable performance in understanding and generating texts. LLMs have large capacity and are trained on vast amounts of data, enabling them to capture complex patterns in natural language with a long context. They have rapidly become powerful tools for various natural language processing tasks. For the QoS prediction problem, each user and service has some specific attributes. For instance, users may have attributes such as its country of origin and the autonomous system it belongs to. Given that LLMs are capable of handling natural language with contextual information, we claim that such capability can be used for QoS prediction task as well. In this paper, we propose to use natural language sentence to describe each user or service using their corresponding attributes. Those attributes contain information such as the proximity and type, which are potentially useful for QoS prediction. We then use state-of-the-art LLMs to extract the descriptive features of users and services from those sentences. This LLMs-based feature helps the QoS prediction model better understand and represent the latent features of users and services. It effectively mitigates the data sparsity issue, as the descriptive LLM features does not rely on historical user-service interactions. It helps the model to provide more reliable and effective service recommendations.
We summarize our main contributions as the following:
  • We first introduce large language models for the web service recommendation task. We propose large language model aided QoS prediction (llmQoS), a novel approach that combines collaborative filtering and nature language processing, making use of both historical user-service interaction information and rich text feature from attributes of users and services. It can effectively mitigate the data sparsity issue inherent to the QoS prediction problem.
  • We show that llmQoS can predict QoS values accurately under different data sparsity levels. The llmQoS model is robust, and can work with different LLMs and even collaborative filtering network architectures. On the WSDream dataset, it outperforms several existing QoS prediction models consistently.
II Related Work
II-A QoS Prediction for Service Recommendation
Among QoS prediction methods for web services, collaborative filtering (CF) [4] methods receive extensive attention and are widely regarded as the most effective approach. CF-based QoS prediction methods can be divided into two main categories: memory-based methods and model-based methods.
Memory-based CF methods first calculate the similarity of users or services, and then predict the QoS values of the target user and service based on historical QoS values and the similarity relationships with their neighbors. Shao et al. [5] first utilize user-based CF methods for QoS prediction. Qi et al. [6] identify “enemies” of the target user to determine their “potential friends”, successfully overcoming the cold-start problem. Meanwhile, item-based methods are widely employed [7, 8]. Chen et al. [9] propose a predictive method that integrates service similarity and geographical location information, effectively alleviating data sparsity and cold-start issues. Additionally, hybrid algorithms that compute both service-to-service and user-to-user similarities [10] are developed. Zheng et al. [11] simultaneously consider user and service similarities for QoS prediction, demonstrating superior predictive performance compared to previous methods.
Model-based CF methods typically use models to predict the QoS values, with matrix factorization (MF) being the most common approach. MF methods decompose the user-service interaction matrix into user and service matrices, predicting values by multiplying these two matrices. Zhu et al. [12] notably extend the traditional MF models by incorporating techniques including data transformation, online learning, and adaptive weighting on the basis of CF techniques. Zheng et al. [13] develop a neighborhood-integrated MF method for personalized web service QoS prediction. Xu et al. [14] employ a probabilistic matrix factorization (PMF) model to learn and predict QoS values, incorporating geographical location to identify user neighbors and comprehensively consider their impact on web service invocation experiences.
Deep learning technology has demonstrated remarkable capabilities across various fields of artificial intelligence in recent years. And it is gradually gaining more attention in the field of QoS prediction too. An increasing number of deep learning based techniques for QoS prediction are developed [15, 16]. Zou et al. [17] propose a novel adaptive QoS prediction framework, utilizing a dual-tower deep residual network to extract latent features of users and services and predict QoS values. Wu et al. [18] introduce a reputation-integrated graph convolutional network for robust and accurate QoS prediction. Zhu et al. [19] develop a twin graph network based on graph contrastive learning, effectively addressing cold-start and data sparsity issues.
Most deep learning based QoS prediction models use the user and service ID information through embedding layers. Additionally, some methods [17, 19] utilize more information of the users and services such as their location, IP address, and the autonomous systems they belong to. Such information provides similarity or proximity signal for the model, which are useful for QoS prediction. In our work, we utilize such information through large language models.
II-B Large Language Model
In the field of natural language processing (NLP), large language models (LLMs) have seen rapid development since the introduction of the transformer architecture [20]. They have shown state-of-the-art performance across virtually all NLP tasks, including text classification and summarizing, machine translation, text generation, and conversation systems.
Most LLMs are based on variants of the already ubiquitous transformer architecture, with the core components being the attention based mechanism that can effectively capture the correlation among different words in the text. It is common for a group of researchers to develop a series of LLMs that share similar architecture and training recipes, but with generational improvement in performance and capability. Notable ones include the BERT series of original BERT[21], RoBERTa [22], ALBERT [23], and DeBERTa [22]; the OpenAI GPT series of GPT [24], GPT-2 [25], GPT-3 [26], and GPT-4 [27]; the META LLaMA series of LLaMA [28], LLaMA2 [29], and LLaMA3 [30]; and the Microsoft Phi series of Phi1 [31, 32], Phi2 [33], and Phi3 [34]. The common trend in each series is that the newer models usually have more parameters, and are trained on larger corpus of text. Larger LLMs have higher capacity, can learn more information from a larger corpus of text, have a higher generalization capability, and in general perform better on standardized NLP benchmarks [35].
Thought LLMs are trained from textual data, they can generalize to tasks beyond NLP too. They have been shown to work on visual question-answering [36], open-vocabulary object detection [37], planning [38], and autonomous driving [39]. In this paper, we use LLMs for the service recommendation task. When using LLMs on non-NLP tasks, it is important to convert the data into a form of prompt that can be processed by the LLMs. In our work we construct descriptive sentences from the relative attributes of web users and services to be used by LLMs. More details are discussed in § III-A.
Different LLMs have different levels of accessibility. The parameters of some LLMs can be downloaded and ran on local computers for any end user, which is often referred as “open-weight” models. Some models are only be used through API calls and the model parameters cannot be accessed. In this paper we only use open-weight models for reproducibility. However our method can be applied to any LLMs with feature extraction functionality.
III Methods
III-A Large Language Model Feature Extraction
FIGURE II: The proposed llmQoS model uses large language models to extract text features from descriptive sentences of user or service. The descriptive sentence is constructed from certain attributes for each user or service, then converted to a sentences-level feature vector, which is used for QoS prediction. Please refer to § III-A for more details.
[…]
After obtaining the descriptive sentence for each user or service, we use pre-trained RoBERTa [22] or Phi3mini [34] to extract sentence-level language features. The pipeline is illustrated in Figure II. The sentence is first converted into a sequence of tokens by a tokenizer. The embedding vector of each token is obtained by their token ID and position. A set of stacked LLM layers process on this sequence of embedding vectors to obtain the final sequence of feature vectors. One of the vector in the final sequence is used as the LLM feature vector of the descriptive sentence.
RoBERTa and Phi3mini have similar overall architecture. Each LLM layer in them comparably consists of an attention based layer [20] and a set of feed-forward layers. The attention layer calculates the correlation among the feature vectors in the sequence. The feed-forward layers are combination of linear layers, normalization layers, and activation functions.
But RoBERTa and Phi3mini have different number of layers, different dimensionality of the embedding vectors, and different tokenizer. Phi3mini has $3.5\times 10^{9}$ parameters, which is significantly larger than RoBERTa’s $1.2\times 10^{8}$ million parameters. They are also pre-trained on different datasets using different training tasks. RoBERTa is pre-trained to do text classification on total of 160 giga-bytes of text data. If each token uses 5 bytes on average, the model sees $3.2\times 10^{10}$ tokens during training. For text classification, the first vector in the final sequence is used. So we also use the first vector as the sentence-level feature. On the other hand, Phi3mini is pre-trained using next token prediction on a dataset with $3.3\times 10^{12}$ tokens. Next token prediction uses the last vector in the final sequence, so we also use the last vector in the final sequence for Phi3mini. Later we show in experiments that Phi3mini outperforms RoBERTa consistently.
III-B Model Architecture
FIGURE III: The overall architecture of the proposed llmQoS model. The embedding features for user and service are combined with the descriptive LLM features described in § III-A, and used by an MLP network and a prediction layer to get the final prediction of the QoS value. Please refer to § III-B for more details.
The overall architecture of the proposed large language model aided QoS prediction (llmQoS) model is illustrated in Figure III. First, both the user and service IDs are projected to feature vectors by their corresponding embedding layers. Formally for user $u$ and service $s$,
$\mathbf{e}_{u}=Embed_{u}(i_{u}),$ (1)
$\mathbf{e}_{s}=Embed_{s}(i_{s}),$ (2)
where $i_{u}$ and $i_{s}$ are the IDs of the user and the service, and $Embed_{u}$ and $Embed_{s}$ are the corresponding embedding layers.
Then descriptive sentences are constructed and the LLM features are extracted for each user and service based on their attributes, as described in § III-A. The descriptive LLM features have much larger dimensionality compared to the ID embedding features. So a projection layer is used for to reduce them to the same dimensionality as the ID embedding features:
$\mathbf{f}_{u}=Proj(LLM(t_{u})),$ (3)
$\mathbf{f}_{s}=Proj(LLM(t_{s})),$ (4)
where $t_{u}$ and $t_{s}$ are the descriptive sentences of the user and service. Please note that we use the same $LLM$ model and projection layer $Proj$ for both user and service.
Then the user ID embedding feature vector, the projected user descriptive LLM feature vector, the service ID embedding feature vector, and the projected service descriptive LLM feature vector are all concatenated. This concatenated feature vector is used as the input to an multi-layer perceptron (MLP) network. Finally a linear layer produces the predicted QoS value for the input user-service pair. Formally,
$\hat{y}=Linear(MLP([\mathbf{e}_{u};\mathbf{f}_{u};\mathbf{e}_{s};\mathbf{f}_{s}])),$ (5)
where $\hat{y}$ is the predicted QoS value for $u$ and $s$.
Please note that our network design is straightforward without complicated architectures such as graph convolution or attention. The QoS prediction network architecture is not the focus of this paper. With straightforward architecture, llmQoS can still achieve state-of-the-art performance as shown by the experiments. This demonstrate the effectiveness of incorporating LLM features for QoS prediction. In § IV-G we further demonstrate that LLM features can be integrated into more complex QoS prediction models and can still increase the performance significantly and consistently.
IV Experiments
TABLE I: The QoS prediction performance in terms of Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of different baseline models and the proposed llmQoS model. Phi3mini is used for llmQoS in this experiment. The best baseline method is marked by gray background, and we show the percentage improvement from the best baseline achieved by llmQoS (Gain). Our proposed model outperforms all baselines at all densities consistently.
Model Throughput
[…]
MAE RMSE MAE RMSE MAE RMSE MAE RMSE
UIPCC 20.757 60.799 22.370 54.456 20.219 50.704 18.928 48.295
RegionKNN 25.632 67.868 24.838 67.551 24.584 67.314 24.036 66.176
LACF 23.169 58.967 19.626 53.105 17.795 49.766 16.667 47.625
PMF 19.082 57.883 15.994 48.071 14.670 44.013 13.924 41.714
BGCL 20.655 61.297 19.318 59.134 18.134 58.804 18.017 58.689
LMF-PP 18.301 51.777 15.913 46.142 14.745 42.993 14.103 41.408
DCALF 17.658 51.632 15.360 46.428 14.384 43.402 13.670 41.624
llmQoS 13.714 46.635 12.022 42.947 11.156 39.567 10.760 38.365
Gain 22.33% 9.68% 21.73% 6.92% 28.93% 7.97% 21.28% 7.35%
Model Response time
[…]
MAE RMSE MAE RMSE MAE RMSE MAE RMSE
UIPCC 0.625 1.388 0.582 1.330 0.501 1.250 0.450 1.197
RegionKNN 0.588 1.543 0.548 1.513 0.526 1.513 0.516 1.521
LACF 0.637 1.444 0.566 1.342 0.516 1.276 0.483 1.230
PMF 0.569 1.537 0.487 1.316 0.452 1.221 0.431 1.169
BGCL 0.461 1.407 0.433 1.374 0.424 1.334 0.416 1.320
LMF-PP 0.529 1.341 0.473 1.242 0.447 1.210 0.426 1.161
DCALF 0.546 1.402 0.486 1.265 0.464 1.210 0.452 1.176
llmQoS 0.409 1.290 0.360 1.224 0.344 1.186 0.327 1.159
Gain 11.36% 3.84% 16.72% 1.41% 18.93% 2.01% 21.40% 0.21%
(a) Throughput $D$=5%
(b) Throughput $D$=10%
(c) Throughput $D$=15%
(d) Throughput $D$=20%
(e) Response time $D$=5%
(f) Response time $D$=10%
(g) Response time $D$=15%
(h) Response time $D$=20%
FIGURE IV: Curves of loss, MAE, and RMSE during the training of llmQoS. The loss continues decreasing while MAE and RMSE on testing set reach minimum earlier. The training epoch when the model achieves the lowest MAE is used as the final model.
IV-A Dataset
To evaluate the performance of our model, we conducted extensive experiments on the WSDream dataset [40]. WSDream is a large-scale real-world dataset of Internet user-service interactions that involves 5,825 services and 339 users in total. Each user and service is identified by a unique ID number. In addition to the IDs, WSDream also provides basic attribute information for each user and service. User attributes consist of IP Address, Country, IP Number, Autonomous System, Latitude, and Longitude. Service attributes include WSDL Address, Service Provider, IP Address, Country, IP Number, Autonomous System, Latitude, and Longitude. We utilize both the IDs and attributes for the users and services in our llmQoS model.
WSDream contains QoS record that include 1,873,850 response time values and 1,831,265 throughput values. For experiments, both the throughput sub-dataset and the response time sub-dataset are randomly split into training and testing sets. We use different density values $D$, which is the percentage of training set data in the total dataset, to test the models’ prediction ability at different data sparsity levels. Specifically we use $D$=5%, 10%, 15%, and 20%. It is expected that with all other settings unchanged, QoS prediction models achieve higher precision with larger values of $D$.
IV-B Implementation Details
[…]
For all experiments, we use Huber loss [43] to train the model. The loss function is defined as
$\mathcal{L}(y,\hat{y})=\begin{cases}\frac{1}{2}(y-\hat{y})^{2},&\text{if }|y-\hat{y}|<\delta\text{;}\\ \delta(|y-\hat{y}|-\frac{\delta}{2}),&\text{otherwise},\end{cases}$ (6)
where $y$ is the ground truth QoS value, and $\hat{y}$ is the QoS value predicted by the model. We use threshold $\delta=1.0$. All models are trained using the Adam optimizer [44] with learning rate of $10^{-4}$ and batch size of 256.
We measure the Mean Average Error (MAE) and Root Mean Squared Error (RMSE) on the testing set to measure the performance of the model. They are defined as
$\text{MAE}=\frac{\sum_{i=1}^{N}|y_{i}-\hat{y}_{i}|}{N},$ (7)
and
$\text{RMSE}=\sqrt{\frac{\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}{N}},$ (8)
where $(y_{i},\hat{y}_{i})$ is a pair of ground truth and predicted QoS values, and $N$ is the number of data points in the testing set. We measure the MAE of the model at the end of each training epoch, and use the one with lowest MAE as the final trained model. We observed that for throughput experiments, the final model can be obtained before 1,500 epochs. For response time experiments, the final model can be obtained before 600 epochs.
IV-C Baseline Models
In our experiment, we selected 7 baselines to compare with our model and evaluate its performance. This includes traditional algorithms and the latest research. We keep dataset split method and evaluation settings for all baselines same to the proposed llmQoS for fair comparison.
UIPCC [11] is a traditional domain-based collaborative filtering method that predicts missing QoS values by leveraging historical invocation information of users and services.
RegionKNN [45] is a hybrid collaborative filtering algorithm that considers geographic information for service recommendations.
LACF [46] is a location-aware collaborative filtering method that integrates user and service location information to recommend web services to users.
PMF [47] is an enhanced traditional matrix factorization method that incorporates probabilistic factors into matrix decomposition for QoS prediction.
BGCL [19] is a graph-based contrastive learning method that utilizes graph attention mechanisms to incorporate domain information of users and services for QoS prediction.
LMF-PP [48] is a location-based matrix factorization method that integrates invocation and domain information, and utilizes fused similarity for preference propagation, addressing cold-start issues to achieve QoS prediction.
DCALF [49] detects domains and noise of users and services from the original QoS matrix, introducing density peak-based clustering methods during modeling for simultaneous detection of neighborhood and noise, achieving QoS prediction.
(a) Throughput $D$=5%
(b) Throughput $D$=10%
(c) Throughput $D$=15%
(d) Throughput $D$=20%
(e) Response time $D$=5%
(f) Response time $D$=10%
(g) Response time $D$=15%
(h) Response time $D$=20%
FIGURE V: QoS prediction performance in terms of Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of llmQoS with different LLM feature extractors. We compare Phi3mini with RoBERTa. Additionally a model without any LLM feature (ID only) is compared. Both LLM based methods outperform ID only model, while larger Phi3mini achieves better QoS prediction performance than RoBERTa.
(a) Throughput $D$=5%
(b) Throughput $D$=10%
(c) Throughput $D$=15%
(d) Throughput $D$=20%
(e) Response time $D$=5%
(f) Response time $D$=10%
(g) Response time $D$=15%
(h) Response time $D$=20%
FIGURE VI: QoS prediction performance in terms of Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of BGCL model with different LLM feature extractors. We compare Phi3mini, RoBERTa, and vanilla BGCL. Similar to Figure V, LLM features are beneficial for QoS prediction, and while larger Phi3mini achieves better QoS prediction performance than RoBERTa.
IV-D Comparison with Baselines
The performance of the proposed llmQoS is compared with the baseline methods in Table I. Phi3mini is used for llmQoS in this experiment. We compare both MAE and RMSE for throughput and response time at different densities. It is evident that llmQoS outperforms all baselines consistently at all densities. For throughput value prediction, llmQoS can reduce the MAE by more than 20%, and RMSE by more than 6% from the best baseline model. For response time value prediction, llmQoS can also reduce MAE for more than 10% from the best baseline method.
The results demonstrate that with the rich features from the descriptive sentences extracted by LLMs, the QoS prediction model can effectively find the similarity and correlation among users and services. Along with the information learned from historical interactions, llmQoS can achieve high QoS prediction accuracy.
IV-E Training Curves
When setting the training epochs, we configured 1500 epochs for throughput prediction and 600 epochs for response time prediction. We plotted training curves to validate these settings. Figure IV illustrates the loss, MAE, and RMSE curves of the llmQoS model during training at sparse densities ranging from 5% to 20%. As the number of iterations increases, the model gradually converges to optimal performance. We observe that the model reaches optimal performance for throughput prediction before 1500 epochs, while for response time prediction, it reaches optimal performance before 600 epochs.
IV-F Comparison of Different LLMs
As discussed in § II-B and § III-A, larger LLMs have higher capacity and can potentially achieve higher performance on various tasks. So here we compare the performance of llmQoS with different LLMs as the feature extractor. We compare the larger and more recent Phi3mini with the smaller and older RoBERTa model. In addition, we remove all LLM feature related components from llmQoS, and only use the ID embedding vectors of users and services to train a QoS prediction model. This model is referred as “ID only” model. We train all models using the same training hyper-parameters, and the results are compared in Figure V.
It is clear that adding LLM features to QoS prediction is helpful, and can reduce the prediction error significantly from the ID only baseline. This further verifies that the information from the attributes of users and services is useful for service recommendation. And Phi3mini achieves lower MAE and RMSE than RoBERTa consistently on all densities. This shows that the superior capability of larger LLMs on NLP tasks also transfers to QoS prediction task.
However, it is important to note that larger LLMs runs at a higher computational cost. Phi3mini is considered one of the smallest modern LLMs, as the larger ones have tens or hundreds of billions of parameters. Larger LLMs can also have higher feature dimensionality, making training the QoS prediction network slower. In practice, trade-off needs to be made between prediction accuracy and computational cost.
IV-G Adding LLM Feature to BGCL
To if LLM features can benefit QoS prediction in general, we further conduct experiments by incorporating LLM features into another QoS prediction model. We selected the Bi-subgraph network based on graph contrastive learning (BGCL) model [19], which has a more complex network architecture. BGCL uses location attributes of users and services for graph contrastive learning, as well as attention mechanism to aggregate features. If LLM features can improve the QoS prediction accuracy for BGCL, we claim that the benefits of using LLM can be further demonstrated.
We keep all other features and network architecture unchanged for BGCL, with the only addition of the LLM feature vectors along side with the embedding vectors of BGCL at the very beginning of the network of BGCL. Similar to § IV-F, we compare BGCL with Phi3mini, BGCL with RoBERTa, and vanilla BGCL. The experimental results, as shown in Figure VI, are similar to the results in Figure V too. We further verify that that integrating LLM features significantly enhances the model’s performance, and larger model achieves more satisfactory results. This demonstrates that LLM features are beneficial to QoS prediction problem, regardless of the network architecture. And we expect this to apply to other QoS prediction models as well.
IV-H Hyper-Parameter Tuning of MLP
(a) Throughput / depth
(b) Response / depth
(c) Throughput / width
(d) Response time / width
FIGURE VII: QoS prediction performance in terms Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of the proposed llmQoS model different depth and width configurations for the MLP network (§ III-B). Depth refers to the number of layers in the MLP, the main experiments use 3 layers. For different width configurations, we multiply the dimensionalities of each layers used by the main experiments by different width factor for comparison. We calculate and show the relative change of MAE and RMSE of a depth or width configuration from the main experiments. This experiment is conducted with density level $D$=5and using Phi3mini. Generally, the effect of configuration of MLP has on the QoS prediction performance is insignificant, with changes less than 5% in most cases.
We test the performance of llmQoS using different configurations for the MLP model as described in § III-B. For this experiment we test the density level $D$=5%, and using Phi3mini as the LLM feature extractor. We vary the depth of the MLP by adding or removing layers from it. The QoS prediction performance of different configurations is shown in 7(a) and 7(b). We change the width of the MLP by multiplying a width factor to the dimensionalities of the layers in it. The QoS prediction performance of different configurations is shown in 7(c) and 7(d).
It can be observed that the depth and width of the MLP network have a minor impact on the QoS prediction performance. In most cases the change in MAE and RMSE is under 5%. However, when the MLP depth is set to 1, or the width factor is set to 0.5, there is a significant deterioration in the performance. This is caused by the network is too shallow and narrow, resulting to very low capacity. Considering this paper primarily focuses on using large language models to assist QoS prediction, we have selected a reasonably appropriate depth and width for the MLP network without extensive searching.
IV-I Training Parameter Tuning
(a) Throughput / BS
(b) Response time / BS
(c) Throughput / LR
(d) Response time / LR
FIGURE VIII: QoS prediction performance in terms Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of the proposed llmQoS model trained using different batch sizes (BS) and learning rates (LR). We calculate and show the relative change of MAE and RMSE of a batch size or learning rate from the main experiments. This experiment is conducted with density level $D$=5and using Phi3mini. The performance is generally insensitive to batch size. It is sensitive to learning rate, with the optimal value around $10^{-4}$.
We test the performance of llmQoS trained using different batch sizes and learning rates. For this experiment we test the density level $D$=5%, and using Phi3mini as the LLM feature extractor. The same configuration of MLP used in the main experiments is used.
The QoS performance of models trained using different batch sizes is compared in 8(a) and 8(b). Generally speaking, the impact of batch size is inconsistent and insignificant. Only for throughput, larger batch size of 1024 causes the prediction accuracy to drop significantly. The performance with different learning rates is compared in 8(c) and 8(d). Learning rate has a huge impact on the model’s prediction accuracy. Very small learning rate results in slow convergence, while large learning rate causes unstable training. Both causes low performance. The optimal learning rate is around $10^{-4}$ as used in the main experiments.
V Conclusion
In this paper, we propose large language model aided QoS prediction (llmQoS) model. It utilizes LLMs to extract features from descriptive sentences constructed based on attributes of web users and services. Then the LLM features are combined with ID based embedding features for QoS prediction through a multi-layer perceptron network. Despite its concise network architecture, llmQoS demonstrates the ability of accurate QoS prediction under data sparsity condition, and outperforms comparable baseline models consistently on the WSDream dataset. We also show that the benefits of adding LLM features is general, and larger LLMs achieves better performance. In the era of explosive development of LLMs, we hope our paper can inspire more researchers in the field of service recommendation to conduct research in this direction.